Two number cubes are rolled three times in a row. What is the probability of rolling a sum of 10 all three times?
Answer by stanbon(75887)

Two number cubes are rolled three times in a row. What is the probability of rolling a sum of 10 all three times?
------
Sum of 10: 4/6 ; 5/5 ; 6/4 ----- 3 ways
--------
P(sum of 10 on one roll) = 3/36 = 1/12
----
P(sum of 10 on 3 rolls) (1/12)^3
=========
Cheers,
Stan H.